OpenStudy (angelberina):

A window consisting of a rectangle topped by a semicircle is to have a perimeter 50. Find the radius of the semicircle if the area of the window is to be a maximum

OpenStudy (dumbcow):

 Perimeter equation: \[2r + 2x + \pi r =50\] Area: \[A = 2rx + \frac{\pi r^2}{2}\] We want to find "r" that maximizes Area solve perimeter for x \[x = 25- \frac{\pi+2}{2} r\] \[A = 50r -2r^2 -\frac{\pi}{2} r^2\] set derivative equal to 0 \[\frac{dA}{dr} = 50-4r - \pi r = 0\] \[r = \frac{50}{\pi+4}\]